# Graph Neural Network Training Systems: A Performance Comparison of Full-Graph and Mini-Batch. [Experiment, Analysis & Benchmark]

### v1

@@ 5.2. Accuracy with Sub-Optimal Setting
Our evaluation on GAT and GCN, which is reported on the extended version (Bajaj et al., 2024), leads to similar conclusions. For GAT, largest difference between FG-train and MB-train is $2.85\%$ in favor of MB-search for the ogbn-arxiv dataset and $1.9\%$ in favor of FG-search for the pubmed dataset. For GCN, it is $1.36\%$ in favor of FG-search for the reddit dataset and $5.5\%$ in favor of MB-search for the pubmed dataset.
Takeaway. These results show that given a dataset and GNN model, each training method can outperform the other depending on the choice of the hyperparameters. This corroborates our choice of performing two separate hyperparameter searches, one per training method, for our comparison. It also potentially explains why existing work shows inconsistent evidence about which training method leads to higher accuracy. We hypothesize that the large variations in accuracy shown in Figure 4 is due to the difference in the choice of hyperparameter settings, which can greatly impact the accuracy and performance of GNN training.
Mini-batch training still performs relatively well in settings where full-graph training performs best, but the opposite is not true. This is an indication that mini-batch training is more robust to different hyperparameter settings, potentially making hyperparameter search easier. We will further corroborate this thesis in later experiments.
5.3. Ease of Hyperparameter Search
We finally answer our last question focusing on accuracy and hyperparameter search, which is Q5) How sensitive are the two training methods to changes in the hyperparameters? We evaluate the ease of finding a good hyperparameter setting using the two training approaches by measuring how accuracy varies as we vary the model architecture hyperparameters, i.e., the number of hidden layers and the size of the hidden vertex dimensions.
We trained each model for a maximum of 1000 epochs, stopping earlier if we the validation accuracy doesn’t increase for 50 epochs.
Figure 6 illustrates the best test accuracy when training the GraphSAGE model, top row shows mini-batch training experiments and the bottom row shows synchronous full-graph training. Notably, mini-batch training demonstrates remarkable consistency across varying model architectures. In contrast, full-graph training exhibits a wider range of variations in test accuracy when the architecture changes. We observed similar results for GCN and GAT models, which are shown in the extended version of this paper (Bajaj et al., 2024).
Figure 6 and the other results for GCN and GAT, which are reported in the extended version (Bajaj et al., 2024), clearly show the effect of over-smoothing, which appears in GNNs at increased depths (Rusch et al., 2023). Mini-batch training on GraphSAGE does not show smoothing with up to 8 layers in Figure 6. In a set of separate experiments we found that oversmoothing does occur with mini-batch training, but only after we increase the number of layers to $16$ or higher. Our experiments for GCN and GAT show that oversmoothing occurs at lower depths for mini-batch training but the resulting accuracy drop is still much less than that with full-graph training (see the extended version (Bajaj et al., 2024)).
Takeaway. These results highlight the robustness of mini-batch training to alterations in model architecture. Full-graph training appears to be more sensitive to changes in architecture, so it requires a more careful and potentially expensive hyperparameter search. Whereas mini-batch training is consistently less sensitive to changes of those parameters than full-graph training, across all the GNN models we consider. In addition, mini-batch training is much less prone to the oversmoothing problem (Rusch et al., 2023) that affects deep GNN models with many layers.
@@ 6. Conclusion
This work is motivated by the need to understand the state of the art in GNN training systems, given the absence of consistent empirical evidence comparing full-graph and mini-batch training systems. It first conducted an in-depth literature survey and investigated the reasons for the inconsistent claims and the lack of empirical evidence in the literature. It further provided a thorough empirical evaluation comparing representative systems supporting the two approaches on various GNNs and datasets.
The study reveals that the mini-batch training systems we considered consistently achieve faster convergence time, measured as time-to-accuracy, compared to full-graph training, regardless of the choice of the dataset, GNN models, and hardware settings. The study also shows that using time-to-accuracy as a metric is more relevant than comparing epoch times for systems that have different convergence criteria. Systems of both types can achieve similar accuracy when hyperparameter tuning is performed separately for full-graph and mini-batch training.
Our work stresses the importance of comparing GNN training systems of both classes in future work in this area. Using time-to-accuracy as the main metric for comparison across areas is important, as is performing separate hyperparameter tuning for the two training methods. These findings are relevant to both machine learning practitioners who utilize GNN training systems and systems researchers involved in their design and evaluation.
@@
The results for highest achievable accuracy for GAT and GCN model are shown in Table 11 and Table 12. Given a dataset, we must compare the accuracies obtained for two combinations: (FG-search, FG-train) for full-graph training and (MB-search, MB-train) for mini-batch training. The results are averaged after 5 runs.
Table 13 shows the best architectures obtained for every model-dataset combination.
@@
Figure 7. Ablation study with number of layers and hidden size for GCN model. Legend represents the size of each hidden layer. X-axis shows the number of layers. Y-axis shows the test accuracy. Top row is for mini-batch training and bottom row is for full graph training. We observe that mini-batch training is robust to architectural changes, whereas full-graph training has a lot of variation.
Figure 8. Ablation study with number of layers and hidden size for GAT model. Legend represents the size of each hidden layer. X-axis shows the number of layers. Y-axis shows the test accuracy. Top row is for mini-batch training and bottom row is for full graph training. We observe that mini-batch training is robust to architectural changes, whereas full-graph training has a lot of variation.
Table 14 shows the results of the time to accuracy experiment for GAT model on a single host with NVLink connectivity.
@@
Table 15 shows the results for the scalability experiments for GAT model on PCIe setup and Table 16 shows the results of the scalability experiment on the NVLink setup.

### v3

@@ 5. Accuracy Evaluation
Our evaluation of GAT and GCN is reported on the appendix in table 9 and table 10 which leads to similar conclusions. For GAT, the largest difference between FG-train and MB-train is $2.85\%$ in favor of MB-search for arxiv and $1.9\%$ in favor of FG-search for pubmed. For GCN, it is $1.36\%$ in favor of FG-search for reddit and $5.5\%$ in favor of MB-search for pubmed.
These results motivate the use of separate hyperparameter tuning for the two methods since good hyperparameter settings do not transfer well across methods.
Impact of optimizations on accuracy. Many performance optimizations we consider can result in a different accuracy than the vanilla methods. This section answers the question Q5: What is the impact of the optimizations on accuracy? In this analysis, we use the best-found hyperparameters for each system class: FG-Search and MB-Search models for full-graph and mini-batch training systems, respectively. For full-graph training, we consider the impact of asynchrony (PipeGCN), sampling (BNS-GCN), and message quantization (AdaQP). For mini-batch training, besides Neighborhood Sampling (NS), we also consider the ClusterGCN and GraphSaint sampling algorithms.
The results for GraphSage are shown in Table 8. Among full-graph training systems, our results show that asynchrony and sampling have minimal impact on accuracy compared to FG, within 1% in all the cases we considered. Message quantization has a stronger negative impact only on arxiv.
Mini-batch training shows a very different trend. For most datasets, there is some sampling algorithm that achieves higher accuracy than all full-graph training approaches. The relative performance of each sampling algorithm shows large variations depending on the dataset. Both GraphSaint and ClusterGCN achieve the best and the worst accuracy among all sampling algorithms for some datasets. These results show that the regularization effect of sampling is key to achieving high accuracy, focusing the training on nodes that have the highest influence on each other (Zeng et al., 2019). Neighborhood Sampling (NS) shows a more stable behavior: on all datasets, it achieves similar accuracy as FG. This justifies its popularity as the default sampling method for mini-batch training. We repeated the experiment on GAT and GCN and obtained similar trends, as shown in the appendix in table 18 and table 19.
Takeaway. When comparing vanilla full-graph and mini-batch training, no method consistently guarantees higher accuracy than the other and the gap between the two methods is not larger than 3.6%. Good hyperparameter settings do not perform as well across both full-graph and mini-batch training, motivating the need for separate tuning processes. The algorithmic optimizations for full-graph training we considered achieve very similar accuracies as the vanilla full-graph method in most cases, making them attractive choices given their performance benefits. For mini-batch training, using the right sampling algorithm can achieve higher accuracy than the full-graph training methods for most datasets and models. However, some sampling algorithms show widely different accuracies based on the dataset and sometimes converge to much lower accuracies. This indicates that trying different algorithms during hyperparameter tuning is essential.
Overall, these results challenge the common rationale for using full-graph training over mini-batch training, suggesting that the expected benefits of avoiding sampling and its associated information loss to achieve higher accuracy do not materialize in practice.
6. Cost Analysis
To support and generalize our empirical observations beyond the specific hardware configurations and software implementations we evaluated empirically, in this section we answer the following question Q6: What are the analytical performance costs of vanilla full-graph and mini-batch training? We model the communication and computation costs of training analytically and then evaluate the cost of pre-processing (sampling) in mini-batch training experimentally to complement the analysis.
Communication cost. Our analysis considers a simplified model of GNN training where the training workload is partitioned among a set of workers $W$, each accessing a local memory. The communication cost is the volume of vertex feature data exchanged among the workers. We ignore the communication cost of gradient synchronization since GNN models have relatively few parameters. We measure the communication cost to convergence rather than the cost per epoch to account following the discussion of Section 4.1.
In full-graph training, at each layer, workers must receive the features of all the vertices in other partitions that have neighbors in the local partition. The communication cost is:
$\Gamma_{\textrm{fg}}=n_{f}\sum_{l=1}^{L}\sum_{w\in W}\sum_{v\in R_{w}}|h_{v}^{l}|$
where $n_{f}$ is the number of epochs to convergence, $R_{w}$ the set of vertices that are remote neighbors of vertices in $w$, $h_{v}^{l}$ is the feature vector for vertex $v$ at layer $l$.
In mini-batch training, we model a scenario where the input features cannot fit in the local memory of one worker and are partitioned among workers. At each iteration, each worker must gather the input features of all the vertices at the bottom layer of its micro-batch. The communication cost to convergence is:
$\Gamma_{\textrm{mb}}=n_{m}\sum_{i=1}^{I}\sum_{w\in W}\sum_{v\in(M_{i,w}\setminus P_{w})}|h_{v}^{0}|$
where $n_{m}$ is the number of epochs to convergence using vanilla mini-batch training, $I$ is the number of iterations in an epoch, $M_{i,w}$ is the set of vertices in the micro-batch assigned to worker $w$ at iteration $i$, and $P_{w}$ is the partition of vertices assigned to worker $w$.
We calculate the ratio between $\Gamma_{\textrm{fg}}$ and $\Gamma_{\textrm{mb}}$ by considering the same hyperparameters for both approaches, which are the same we used for evaluating the time-to-accuracy in Section 4.1. We use Metis to partition the datasets (Karypis and Kumar, 1997) and run Neighborhood Sampling for one epoch to obtain $M^{i}_{w}$.
Figure 7 shows the ratio between $\Gamma_{\textrm{fg}}$ and $\Gamma_{\textrm{mb}}$ for different datasets. The communication cost of full-graph training is higher than that of mini-batch training in most cases. The gap mostly depends on the number of boundary nodes and, for mini-batch training, on the size of the last layer of the micro-batches. In practice, mini-batch training does not need to partition the dataset since it can replicate commonly accessed input features in the local memory of multiple workers. In our experimental evaluation of Section 4, all datasets except orkut and papers100M are fully replicated, but partial replication is often almost as effective (Song et al., 2023b).
In general, larger graphs can have a larger edge cut when they are partitioned. This results in larger $R_{w}$, increasing the communication cost of full-graph training $\Gamma_{\textrm{fg}}$. In contrast, the communication cost of mini-batch training, $\Gamma_{\textrm{mb}}$, grows with $(M_{i,w}\setminus P_{w})$, which is upper bounded by the size of the sampled micro-batches, not by the size of the graph. A similar effect is observed by increasing the number of partitions with fixed datasets, as shown in Figure 7. This increases the relative cost of $\Gamma_{\textrm{fg}}$ over $\Gamma_{\textrm{mb}}$.
Figure 7. FG/MB ratio of communication cost. (Left: Hyperparameters in the appendix in table 16. Right: Default 4 partitions)
Figure 8. FG/MB ratio of computation cost for GraphSage.
Computation cost. We now analyze the computational cost of training with vanilla full-graph and mini-batch training. We consider the same model used for the communication cost analysis.
A GNN training layer $l$ computes the features of vertices at layer $l$ based on the features of their neighbors at layer $l-1$ (see Section 2.1). The computational cost to the convergence of full-graph training can be expressed as:
$\Theta_{\textrm{fg}}=n_{f}\sum_{l=1}^{L}(|E^{l}|\cdot c_{e}+|V^{l}|\cdot c_{v})(1+\eta)$
where the summation expresses the cost of the forward pass: $c_{e}$ is the cost of generating and aggregating a message sent over one edge, $c_{v}$ is the cost of computing a new representation of a vertex, $V^{l}$ is the set of all vertices processed at layer $l$ in the epoch, and $E^{l}$ is the set of all edges between $V^{l-1}$ and $V^{l}$ processed in the epoch. The factor $(1+\eta)$ accounts for the cost of the backward pass.
The values of $c_{e}$ and $c_{v}$ are independent of the choice of full-graph or mini-batch training. They depend on the GNN model (e.g. GraphSage or GAT) and its hyperparameters and they can be calculated analytically in terms of FLOPS.
The computational cost for mini-batch training is similar:
$\Theta_{\textrm{mb}}=n_{m}\sum_{i=1}^{I}\sum_{w\in W}\sum_{l=1}^{L}(|E^{l}(M_{i,w})|\cdot c_{e}+|V^{l}(M_{i,w})|\cdot c_{v})(1+\eta)$
where $E^{l}(M_{i,w})$ and $V^{l}(M_{i,w})$ are the number of edges and vertices at layer $l$ in the micro-batch $M_{i,w}$.
We calculate the ratio of $\Theta_{\textrm{fg}}$ and $\Theta_{\textrm{mb}}$ using the same hyperparameters as in the communication cost analysis and show it in Figure 8. The results reveal a large difference in terms of computational cost between mini-batch and full-graph training, especially with the largest graph, papers100M. In full-graph training, $\Theta_{\textrm{fg}}$ depends on $|E^{l}|$ and $|V^{l}|$, which grow as the size of the graph grows. Mini-batch training performs message passing only on a sample of the graph, unlike full-graph training, reducing computation cost substantially. The computation cost $\Theta_{\textrm{mb}}$ is upper-bounded by the size of $M_{i,w}$, which depends on the number of layers in the GNN and the sampling algorithm rather than on the size of the graph.
As the number of workers grows, the gap between full-graph and mini-batch training decreases in Figure 8. This is because the same vertex can appear in multiple micro-batches, so its hidden features are computed multiple times in an epoch. This redundant computation does not happen with full-graph training.
Cost of sampling in mini-batch training. Mini-batch training requires two pre-processing steps: sampling and data loading. Data loading costs are included in the communication costs. Modeling sampling costs in terms of FLOPS, as we did for the training, costs $\Theta_{\textrm{mb}}$ is challenging because sampling is a sparse and irregular computation, not a dense computation like matrix multiplication. Therefore, to put the previous results in context, it is necessary to consider the relative computational cost of sampling over training in mini-batch training. We measure it experimentally using DGL and Neighborhood Sampling, running both sampling and training in CPU memory to factor out the overhead of transferring data to the GPUs. The results are shown in Figure 9. The computational cost of sampling is noticeable but is still lower than the training cost. It decreases when we increase the number of workers/partitions because each epoch requires fewer iterations.
Takeaway. This analysis shows that vanilla full-graph training has a higher communication cost than vanilla mini-batch training, especially with larger graphs. In practice, mini-batch training can replicate input features across multiple workers to eliminate or reduce communication costs. The high communication cost of full-graph training has been the focus of prior work on full-graph training systems, as discussed in Section 2.2.
The analysis shows that using sampling in mini-batch training also results in a substantial reduction in computation cost compared to full-graph training systems, especially with larger graphs. Overcoming this gap while maintaining the full-graph training approach and avoiding sampling is challenging.
Figure 9. Fraction of epoch time spent performing sampling in mini-batch training for GraphSage.
@@ 7. Conclusion
A common rationale for using full-graph training is to achieve high accuracy by avoiding the information loss results from sampling. In our evaluation, we did not observe these expected benefits. Mini-batch training systems not only achieve a lower time-to-accuracy across all models, datasets, and hardware configurations we consider but also a comparable or even higher accuracy with proper hyperparameter tuning.
Mini-batch training also has some limitations. The common Neighborhood Sampling algorithm is able to consistently achieve good accuracy, but it still shows a slightly larger variance between training runs than vanilla full-graph training. Other sampling algorithms we evaluated are not always able to converge to a high accuracy, depending on the dataset. Finding sampling algorithms that have consistently higher performance than Neighborhood Sampling is an interesting avenue of research. Our experiments show that as GNN models become deeper, the advantage of mini-batch training over full-graph training reduces. Distributed sampling over multiple hosts can also become a bottleneck, especially when the network is slow.
An important advantage of mini-batch training systems is that the sampling algorithms are interchangeable components. This modularity and flexibility in choosing the right sampling algorithm is key since different algorithms perform very differently on different datasets. Our experiments show that providing efficient sampling implementations is important for multi-host setups, especially when using a slow network. In order to support future innovations in sampling algorithms, it is important to have systems that allow ML practitioners to implement new sampling algorithms and run them on GPUs, such as (Jangda et al., 2021; Pandey et al., 2020; Gong et al., 2023; Wang et al., 2021).
In terms of improving full-graph training systems, our cost model analysis and single-GPU performance experiments reveal that in addition to the well-known communication cost of full-graph training—a bottleneck addressed by much previous work—there is a computation cost gap that is difficult to bridge without employing sampling. A natural open question is whether combining full-batch training with more aggressive sampling, forgoing full-graph aggregation, can be more effective than standard mini-batch training. In our hyperparameter tuning, however, we found that the accuracy of mini-batching typically peaks with batch sizes around 1024/4096 and does not grow further with larger values, so the advantages of using full-batch training are unclear.
@@
The results for the highest achievable accuracy for the GAT and GCN model are shown in Table 9 and Table 10. Given a dataset, we must compare the accuracies obtained for two combinations: (FG-search, FG-train) for full-graph training and (MB-search, MB-train) for mini-batch training. The results are averaged after 5 runs.
Table 11 shows the search space of our hyperparameter search. Table 12 shows the best architectures obtained for every model-dataset combination. Table 13 shows the architectures used for TTA comparisons. Tables 14 and  15 show the architectures used for scalability comparisons. Table 16 shows the architectures used for communication cost experiments.
@@
Table 11. Hyperparameter search space
Parameter range/values
Number of Hidden Layers [2, 10]
Size of hidden layers [16, 1024]
Aggregator (GraphSage only) { mean, gcn, pool }
Number of heads (GAT only) [1, 12]
Sampling Fanout (MB only) [4, 50]
Learning rate [0.0001, 0.01]
Dropout [0.2, 0.9]
Micro-batch size (MB only) [32, 1024]
@@
Fanout 10 15 15 5
Table 13. Architectures used for TTA experiments (same for FG and MB).
Model Architecture Pubmed Ogbn-arxiv Reddit Ogbn-products Ogbn-papers100M
GraphSAGE Layers 3 2 4 5 2
Hidden size 256 512 1024 256 128
Fanout 10 20 5 5 5
GAT Layers 3 3 2 3 -
Hidden size 256 1024 1024 128 -
Fanout 10 10 10 5 -
Num Heads 4 6 4 3 -
Table 14. Architectures used for Scalability experiments single host (same for FG and MB).
Model Architecture Ogbn-arxiv Reddit Ogbn-products Orkut
GAT Layers 3 2 3 3
Hidden size 1024 1024 128 128
Num Heads 6 4 3 2
Fanout 25 20 20 5
Table 15. Architectures used for Scalability experiments for 3 hosts (same for FG and MB).
Model Architecture Orkut Ogbn-papers100m
GraphSAGE Layers 3 2
Hidden size 512 512
Fanout 10 10
Table 16. Architectures used for Communication cost experiments (Same for FG and MB).
Model Architecture Pubmed Ogbn-arxiv Reddit Ogbn-products Ogbn-papers100M
GraphSAGE Layers 3 2 4 5 2
Hidden size 256 512 1024 256 128
Fanout 10 20 5 5 5
@@
Table 17 shows the results of the time-to-accuracy experiment for the GAT model on a single host with NVLink connectivity. Table 18 and  19 show the accuracies achieved for GAT and GCN models for different training techniques an optimizations.
Table 20 shows the results of the scalability experiments for the GAT model on PCIe setup and table 21 shows the results of the scalability experiment on the NVLink setup. Table 22 shows the results of the time-to-accuracy experiments for the GraphSAGE model on 3 PCIe hosts.
Tables 23, 24 and 25 show the epoch time and time to accuracy for GraphSAGE, GAT and GCN respectively for different sampling optimizations. Figure 13 shows the best accuracy achieved and 14 shows the time-to-accuracy plots for GraphSAGE, GAT and GCN models with different mini-batch optimizations (Neighbor sampler, ClusterGCN and SAINT sampler).
Table 18. Best accuracies (%) for GAT model
Pubmed Ogbn-arxiv Reddit Ogbn-products
FG $77.70\pm 0.50$ $71.74\pm 0.22$ $95.17\pm 0.20$ $74.43\pm 0.05$
PipeGCN $77.60\pm 0.37$ $70.34\pm 0.10$ $93.63\pm 0.07$ $74.08\pm 0.12$
BNS-GCN $77.70\pm 0.10$ $71.75\pm 0.70$ $94.23\pm 0.02$ $74.50\pm 0.26$
NS $78.50\pm 0.03$ $70.63\pm 0.12$ $95.58\pm 0.13$ $74.57\pm 0.07$
ClusterGCN $80.07\pm 0.19$ $63.30\pm 0.09$ $92.49\pm 0.75$ $73.27\pm 0.29$
GraphSaint $82.76\pm 0.95$ $72.58\pm 0.17$ $96.96\pm 0.06$ $70.47\pm 0.16$
Table 19. Best accuracies (%) for GCN model
Pubmed Ogbn-arxiv Reddit Ogbn-products
FG $79.40\pm 0.12$ $71.54\pm 0.21$ $94.67\pm 0.20$ $75.71\pm 0.18$
PipeGCN $77.00\pm 0.81$ $69.64\pm 0.27$ $93.91\pm 0.10$ $74.25\pm 0.34$
BNS-GCN $77.70\pm 0.71$ $70.75\pm 0.36$ $93.23\pm 0.09$ $74.50\pm 0.32$
NS $78.70\pm 0.23$ $71.00\pm 0.16$ $94.41\pm 0.14$ $79.31\pm 0.04$
Cluster $77.50\pm 1.98$ $63.24\pm 0.10$ $92.42\pm 0.03$ $73.71\pm 0.08$
GraphSaint $81.70\pm 0.71$ $69.72\pm 0.08$ $96.88\pm 0.06$ $72.65\pm 0.13$
@@
Table 22. Time to accuracy (TTA) and epoch time (ET) for GraphSage and ogbn-papers100M (3 PCIe hosts)
System Time (s)
Epoch time (s) Full-Graph 22.00
PipeGCN 1.03
BNS-GCN 2.46
DistDGL 38.37
Time to accuracy (s) Full-Graph 19141.48
PipeGCN 956.34
BNS-GCN 859.26
DistDGL 767.38
Appendix A Ease of Hyperparameter Search
Figure 11. Ablation study with varying number of layers and hidden size for GCN model. Legend represents the size of each hidden layer. The X-axis shows the number of layers. The Y-axis shows the test accuracy. The top row is for mini-batch training and the bottom row is for full graph training. We observe that mini-batch training is robust to architectural changes, whereas full-graph training has a lot of variation.
Figure 12. Ablation study with the varying number of layers and hidden size for GAT model. Legend represents the size of each hidden layer. The X-axis shows the number of layers. The Y-axis shows the test accuracy. The top row is for mini-batch training and the bottom row is for full graph training. We observe that mini-batch training is robust to architectural changes, whereas full-graph training has a lot of variation.
How sensitive are the two training methods to changes in the hyperparameters? We evaluate the ease of finding a good hyperparameter setting using the two training approaches by measuring how accuracy varies as we vary the model architecture hyperparameters, i.e., the number of hidden layers and the size of the hidden vertex dimensions.
We trained each model for a maximum of 1000 epochs, stopping earlier if the validation accuracy doesn’t increase for 50 epochs.
Figure 10 illustrates the best test accuracy when training the GraphSAGE model, the top row shows mini-batch training experiments and the bottom row shows synchronous full-graph training. Notably, mini-batch training demonstrates remarkable consistency across varying model architectures. In contrast, full-graph training exhibits a wider range of variations in test accuracy when the architecture changes. We observed similar results for GCN and GAT models, which are reported in Figure 11 and Figure 12.
Figure 10 and the other results for GCN and GAT, clearly show the effect of over-smoothing, which appears in GNNs at increased depths (Rusch et al., 2023). Mini-batch training on GraphSAGE does not show smoothing with up to 8 layers in Figure 10. In a set of separate experiments, we found that over-smoothing does occur with mini-batch training, but only after we increase the number of layers to $16$ or higher. Our experiments for GCN and GAT show that over-smoothing occurs at lower depths for mini-batch training but the resulting accuracy drop is still much less than that with full-graph training (see the Appendix).
Takeaway. These results highlight the robustness of mini-batch training to alterations in model architecture. Full-graph training appears to be more sensitive to changes in architecture, so it requires a more careful and potentially expensive hyperparameter search. Whereas mini-batch training is consistently less sensitive to changes of those parameters than full-graph training, across all the GNN models we consider. In addition, mini-batch training is much less prone to the over-smoothing problem (Rusch et al., 2023) that affects deep GNN models with many layers.
Table 23. Time to accuracy and epoch time for GraphSAGE
Sampler pubmed ogbn-arxiv reddit ogbn-products
Epoch time (s) Cluster GCN 0.0118 0.0254 0.2798 0.1212
NS 0.0120 0.1592 2.6728 0.5552
Saint Sampler 0.0123 0.1029 0.4689 0.0908
Time to accuracy (s) Cluster GCN 2.5418 9.0578 97.6485 111.46
NS 1.1168 48.3875 192.440 122.71
Saint Sampler 1.0939 61.2358 151.910 42.9602
Table 24. Time to accuracy and epoch time for GAT
Sampler pubmed ogbn-arxiv reddit ogbn-products
Epoch time (s) Cluster GCN 0.0353 0.0621 1.3446 0.4893
NS 0.0222 0.2927 18.5336 1.2580
Saint Sampler 0.0224 0.1176 0.3653 0.0524
Time to accuracy (s) Cluster GCN 2.3293 15.3941 599.686 1109.32
NS 3.8463 46.2388 1853.36 159.769
Saint Sampler 2.4866 47.759 115.419 20.4534
Table 25. Time to accuracy and epoch time for GCN
Sampler pubmed ogbn-arxiv reddit ogbn-products
Epoch time (s) Cluster GCN 0.0112 0.0259 0.2333 0.1030
NS 0.0133 0.1814 4.4802 0.6186
Saint Sampler 0.0126 0.1016 0.4506 0.0905
Time to accuracy (s) Cluster GCN 2.5933 9.1563 39.6631 47.0832
NS 1.8819 27.2038 488.345 64.3342
Saint Sampler 2.5587 51.6232 159.962 29.0637
Figure 13. Test accuracy achieved for the best architecture for GraphSAGE, GAT and GCN models with NeighborSampler, ClusterGCN and SAINT Sampler.
Figure 14. Time to accuracy plots for GraphSAGE, GAT and GCN models with NeighborSampler, ClusterGCN and SAINT Sampler.